OpenStudy (anonymous):

Solve the exponential equation. Express the solution set in terms of natural Logarithms. 3^x+8=5

OpenStudy (raden):

3^x + 8 or 3^(x+8) ?

OpenStudy (anonymous):

x+8 is the exponent of 3

OpenStudy (raden):

oh ok, it is 3^(x+8) = 5 put ln on both sides, get ln 3^(x+8) = ln 5 (x+8) ln 3 = ln 5 x+8 = ln5/ln3 x = ln5/ln3 - 8

OpenStudy (anonymous):

Thank you

OpenStudy (raden):

welcome